OpenStudy (anonymous):

what is 4*3/4 = ????? please explain

OpenStudy (anonymous):

is this: \[(4)* (3/4)\] or \[(4*3)/(4)\]

OpenStudy (anonymous):

More likely the latter is the case. PEMDAS right?

OpenStudy (anonymous):

thats too easy then

OpenStudy (anonymous):

(4*3)/(4)=12/4=3

OpenStudy (anonymous):

This is like that Question that raged across Facebook >.> And haha Lagrange

jimthompson5910 (jim_thompson5910):

btw LagrangeSon, \[(4)\left(\frac{3}{4}\right)=\frac{(4*3)}{4}\] So in this case, it really doesn't matter.